OpenStudy (anonymous):

-3x + 4y = 80 x + 4y = 48 What are the values of x and y?

OpenStudy (uri):

-3x + 4y = 80 x + 4y = 48 - - - ======== -4x=32 x=-32/4 x=-8 -8+4y=48 4y=48+8 4y=56 y=56/4 y=14 x=-8 y=14

OpenStudy (anonymous):

Thank you. I was completely lost. I'm not the best in math lol.